Mrs. Lopez is 54 and has a daughter who is six years more than a third of her age. How old is Mrs. Lopez's daughter?

Answer by omdenman(1)

The daughter is 24 years old. To find one third of Mrs. Lopez's age you do 54/3=18. Then you add 6 more to it. 18+6=24
54/3+18
18+6=24